# Let’s Fuse Step by Step: A Generative Fusion Decoding Algorithm with LLMs for Multi-modal Text Recognition

### v3

@@ Abstract
We introduce “Generative Fusion Decoding” (GFD), a novel shallow fusion framework, utilized to integrate Large Language Models (LLMs) into multi-modal text recognition systems such as automatic speech recognition (ASR) and optical character recognition (OCR). We derive the formulas necessary to enable GFD to operate across mismatched token spaces of different models by mapping text token space to byte token space, enabling seamless fusion during the decoding process. The framework is plug-and-play, compatible with various auto-regressive models, and does not require re-training for feature alignment, thus overcoming limitations of previous fusion techniques. We highlight three main advantages of GFD: First, by simplifying the complexity of aligning different model sample spaces, GFD allows LLMs to correct errors in tandem with the recognition model, reducing computation latencies. Second, the in-context learning ability of LLMs is fully capitalized by GFD, increasing robustness in long-form speech recognition and instruction aware speech recognition. Third, GFD enables fusing recognition models deficient in Chinese text recognition with LLMs extensively trained on Chinese. Our evaluation demonstrates that GFD significantly improves performance in ASR and OCR tasks, with ASR reaching state-of-the-art in the NTUML2021 benchmark. GFD provides a significant step forward in model integration, offering a unified solution that could be widely applicable to leveraging existing pre-trained models through step by step fusion.  [^1] Code is available at https://github.com/mtkresearch/generative-fusion-decoding
@@ 1 Introduction
In recent years, integrating Large Language Models (LLMs) into multi-modal systems has emerged as a frontier, offering substantial advancements in applications such as automatic speech recognition (ASR) [40], visual question answering (VQA) [36], reinforcement learning [50], and generalist agent [42]. The robust language model capabilities that the LLM offers make it an ideal candidate for integration with other models, many of which are “undertrained” in the textual domain, compared to trillions of tokens consumed by the LLMs. Obtaining sufficient quality paired data and training resources in a huge endeavor for text recognition systems such as automatic speech recognition (ASR) and optical character recognition (OCR).
The integration of heterogeneous models has been explored extensively in ASR literature, as exemplified by strategies such as shallow fusion, late fusion, mid fusion, and early fusion. Works in shallow fusion [9, 27, 10] interpolates text recognition outputs with LM outputs at each step of beam search. Late fusion approaches typically engage LLMs to score the initial sentence generated by the recognition model. The scores are then utilized for fusing with recognition probabilities for joint transcription [6], for re-ranking the the constrained set of hypothesis as work in re-scoring [46], for generating new hypothesis options through Generative Error Correction (GER) [5]. Notably, methods in mid-fusion propose fusing models through model attentions to integrate LLM into the decoding process of the recognition model [41, 37], and work in early-fusion has proposed concatenating speech embeddings or image embeddings along with LLM embedings as inputs to the LLM for ASR and VQA [17, 7]. However, the aforementioned fusion methods bear several challenges. The early fusion techniques are often limited by a low correction rate due to the restricted range of available choices for re-evaluation. Mid-fusion discards the whisper decoder, which has been shown to have denoising ability[19]. Additionally, late fusion with GER may lead to hallucination due to the LLM as the LLM generation is devoid of auditory inputs. Due to the requirements for shallow and late fusion to align output spaces, and the need for mid and early fusions to involve joint training with additional paired data, prevent an LLM from being used as is with existing text-recognition models without proper feature alignment through trainings.
Figure 1: The GFD integrated framework. The framework aims to integrate pre-trained text-recognition models with LLMs to augment the recognition capabilities. However, direct fusion is impeded by the discrepancy in token spaces between the two model types. To overcome this obstacle, GFD is introduced, which converts the token space into a byte space, thereby facilitating incremental fusion. Equations 7 and 8 are employed to compute the fusion probability, enabling a stepwise fusion process. The symbols $\mathcal{Z}^{\texttt{TR}}$ and $\mathcal{Z}^{\texttt{LM}}$ denote the contextual information pertinent to the text-recognition models and the language models, respectively. Notably, $\mathcal{Z}^{\texttt{TR}}$ may correspond to either an audio input for ASR or a text image for OCR, whereas $\mathcal{Z}^{\texttt{LM}}$ typically represents a textual prompt.
To address the aforementioned challenges, we introduce a novel shallow fusion framework, termed “Generative Fusion Decoding” (GFD). GFD resolves the token space mismatch between models by bridging heterogeneous spaces through the expected score values of plausible sequences and mapping text token space to byte token space. This allows GFD to seamlessly fuse LLMs with text recognition models during the decoding process. Our unified framework is plug-and-play, applicable to text recognition tasks such as ASR and OCR, and compatible with diverse auto-regressive text recognition models as well as language models. We highlight three main advantages of GFD. First, GFD addresses key limitation of previous fusion techniques — namely the complexity of aligning different model sample spaces via re-training associated with prior deep integration [41] or two-pass [46] strategies. It allows LLMs to participate in the sampling process as conventional shallow fusion, providing real-time correction of errors generated by the text recognition model, thereby extending the influence of LLMs beyond the complete outputs of the recognition model. Secondly, as modern LLMs are trained on trillions of multilingual text tokens [22, 26], generally around 3 orders of magnitude more than end-to-end ASR or OCR models [40], LLMs in GFD can be seen as a strong scoring model in the step by step decoding that improves the multi-lingual outputs of ASR and OCR systems. Lastly, because of the in-context learning ability of LLMs, the GFD fusion system with LLMs and ASR can readily leverage domain information as prompts effectively for domain adaptation similarly to prior work through re-scoring [14] or instruction tuning [35].
To evaluate the efficacy of our algorithm, we implemented the GFD for ASR and OCR tasks. In the ASR task, we evaluated on the conventional ASR task as shown in Section 4.2 for English and Chinese audios. The evaluation of Chinese ASR is motivated by existing literature [40], which indicates a performance outlier in Chinese compared to other languages despite abundant paired-training data. We hypothesized that LLMs could enhance Chinese ASR by offering additional scoring to differentiate between semantically distinct but homophonous words. Furthermore, we demonstrate that ASR with GFD can improve recognition for long-form audio through leveraging the long-context ability of LLMs (Section 4.3); for context-aware recognition for rare words and domain tags via in-context learning capabilities of LLMs (Section 4.4). Lastly, application to the OCR task, particularly targeting long text OCR for historical documents, demonstrates that OCR with GFD can significantly improve recognition accuracy (Section 4.5).
@@ 2.1 Model fusion
Training a multi-objective model from scratch is often costly [3, 2, 1, 15]. Consequently, researchers have pivoted towards combining existing models with different modalities to improve accuracy without the prohibitive costs of building new systems from the ground up.
Model fusion developed in the field of ASR provides a plausible path for combining existing trained models. The technique have evolved significantly in recent years, encompassing a variety of approaches designed to integrate different models to enhance performance. Shallow fusion or late fusion, often employed in ASR, combines end-to-end ASR models with external language models at the decoding level, improving recognition accuracy without altering the underlying ASR architecture [27, 25, 6, 52]. However, due to the heterogeneous sample spaces of models, the prerequisite of shallow and late fusion requires aligning sample spaces of model distributions, enabled through fine-tuning a projection module [6]. Late fusion training methods may suffer from modality laziness problem in tasks where uni-modal priors are meaningful [16]. Work in re-scoring N-best hypothesis from the first pass inference with an external language model has been shown effective in reducing recognition errors [43, 23, 46]. However, the method prunes the correct prefix early on away from N-best list and requires a second pass in the decoding process.
Alternatively, deep fusion integrates models at the level of hidden feature, requiring fine-tuning models to fuse deep features [20]. Cross-modal fusion, similar to deep fusion, integrates pre-trained end-to-end ASR model with LLM [41, 51, 33] or vision model with LLM [7, 36] via learning a joint representation with large amount of extra paired audio-text or image-text data.
Our newly proposed approach—GFD—decodes in homogeneous sequence element space with byte strings, and therefore relaxes the alignment of sample spaces of distributions without any model training. Without pre-training or fine-tuning, we ensure that the individual pre-trained model performance is guaranteed on their respective task, and is not influenced by volatility of further training.
@@ 2.2 Contextual conditioning
Auto-regressive LLMs have exhibited capabilities in in-context learning [39], instruction following [38], and knowledge synthesis [44]. Such capabilities have been applied to solving domain adaptation in speech recognition for rare words or out-of-domain context through contextual biasing [10] and prompting fine-tuned models [35, 47, 33, 47]. With GFD, such problem setting can be tackled by leveraging a performant LLM as is through prompting without further fine-tuning.
2.3 Chinese ASR
One of the most significant challenges in developing ASR systems for the Chinese language stems from its highly homophonous nature [30, 29, 8]. Unlike English, where there is a larger variety of phonemes and a relatively consistent correspondence between spelling and sound, Chinese relies on a limited set of tones and syllables to represent thousands of characters. Consequently, Chinese ASR systems must not only accurately capture the tonal nuances but also analyze the linguistic context to disambiguate these homophones. This requires sophisticated LMs that can effectively leverage contextual clues to predict the intended characters. The integration of LLMs has shown promise in addressing these challenges [11, 31, 34]. GFD is also expected to enhance the performance of Chinese ASR by utilizing the contextual conditioning capabilities of LLMs.
@@ 3.1 Generative fusion decoding
where $\{T_{s}\}$ represents the sequence of tokens generated by the model, and $\mathcal{Z}$ represents the given context or conditioning information, such as audio for speech recognition models [40], images for vision-language-models [1], and prompts for typical language models [4].
Auto-regressive generation is one approach to realize conditional text generation. In this approach, the auto-regressive model is conditioned on the previously generated tokens to generate the next token sequentially. Therefore, the probability $\log\mathbb{P}(\{T_{s}\}|\mathcal{Z})$ is typically decomposed using the chain rule of probability, as follows:
@@
where $T_{s}$ is the token at position $s$ in the sequence, and $T_{<s}$ represents all the tokens preceding position $s$. In real-world applications, it is impracticable to enumerate all possible token sequences, so beam search is typically used as an approximation method.
In the setting of shallow fusion, we attempt to let multiple models jointly determine the sequence, as shown in the following formula rewrote from Equation (1):
$\{T_{s}^{\texttt{fuse}}\}^{*}$ $=\underset{\{T_{s}^{\texttt{fuse}}\}}{\mathrm{arg\,max}}\sum_{m}\lambda_{m}\log\mathbb{P}_{m}(\{T_{s}^{(m)}\}=\{T_{s}^{\texttt{fuse}}\},\mathcal{Z}^{(m)})$ (3)
where $\{T_{s}^{\texttt{fuse}}\}$ represents the fused sequence of tokens generated by combining the outputs of multiple models, $\lambda_{m}$ is a weighting factor for the $m$-th model, $p_{m}$ denotes the probability distribution of the $m$-th model and $\mathcal{Z}^{(m)}$ represents the context or conditioning information specific to the $m$-th model.
When the sample spaces of models are the same, Equations (3) and (2) can be combined to realize incremental fusion. If the models have different sample spaces due to a mismatch in token spaces, there is no simple way to achieve incremental fusion. One alternative method to approximate Equation (3) is to fuse at the level of the fully generated results from each model. Nevertheless, in practice, fusion at the level of fully generated results poses the problem of an enormous search space because different conditioning variables $\mathcal{Z}^{(m)}$ may produce vastly different results.
@@
$\mathcal{M}^{(m)}:\mathbb{P}_{m}(\{T_{s}\},\mathcal{Z}^{(m)})\longrightarrow\mathbb{P}_{m}(\{B_{l}\},\mathcal{Z}^{(m)}),$ (4)
@@
$\{B_{l}^{\texttt{fuse}}\}^{*}$ $=\underset{\{B_{l}^{\texttt{fuse}}\}}{\mathrm{arg\,max}}\sum_{m}\lambda_{m}\log\mathbb{P}_{m}(\{B_{l}\}=\{B_{l}^{\texttt{fuse}}\},\mathcal{Z}^{(m)}).$ (5)
To operationalize the probability transformation $\mathcal{M}^{(m)}$, we define a mapping from the token-level probabilities to the byte-level probabilities. This mapping takes into account the prefix relationship between the token sequence and the byte sequence. Specifically, we express the byte-level probability $\mathbb{P}_{m}(\{B_{l}\},\mathcal{Z}^{(m)})$ as a sum over all possible token sequences that share a common prefix with the byte sequence $\{B_{l}\}$. The probability of each token sequence is computed as the product of the conditional probabilities of each token given the preceding tokens and the context $\mathcal{Z}^{(m)}$. This relationship is formalized in the following equation:
$\mathbb{P}_{m}(\{B_{l}\},\mathcal{Z}^{(m)})=\sum_{\{B_{l}\}}\biggr{[}\prod_{s}\mathbb{P}_{m}(T_{s}^{(m)}\mid T_{<s}^{(m)},\mathcal{Z}^{(m)})\biggr{]}_{\{T_{s}\}}\times\mathbb{1}\bigl{(}\{T_{s}\}\texttt{.prefix}=\{B_{l}\}\bigl{)},$ (6)
where .prefix is a function that maps a token sequence to its corresponding byte sequence prefix, and $\mathbb{1}$ is the indicator function that is 1 if the token sequence has the specified byte sequence as a prefix and 0 otherwise. The indicator function ensures that only those token sequences that can be mapped to the given byte sequence contribute to the byte-level probability. In Equation (6), the complexity remains high at $\mathcal{O}(V^{S})$, where $V$ represents the vocabulary size of tokens and $S$ is the sequence length, for identifying token sequences that match the specified criteria of the indicator.
@@
To reduce complexity, we posit that the main token sequence and its branching alternatives, as shown in Figure 2, dominate the probability contribution, where the main token sequence is produced by applying model tokenization on the byte string and the alternative tokens are essentially look-aheads for potential main tokens that may emerge as the decoding progresses. The significance of the main token sequence is justified by its alignment with the model’s inputs during the original pretraining phase; any other slicing method is penalized in terms of probability due to its lack of representation in the training data. Based on this assumption, we can narrow down the search for token sequences that meet the criteria of $\{T_{s}\}\texttt{.prefix}=\{B_{l}\}$ to only the main token sequence and its branching alternative tokens. We define the main token sequence as $\{\hat{T}_{s}\}$. Given this simplification, we can approximate the byte-level probability $\mathbb{P}_{m}(\{B_{l}\},\mathcal{Z}^{(m)})$ by considering only the main token sequence and its immediate alternatives that share the same prefix with the byte sequence $\{B_{l}\}$. This approximation significantly reduces the computational complexity to $\mathcal{O}(V\times S)$ and is expressed in the following equation:
$\mathbb{P}_{m}(\{B_{l}\},\mathcal{Z}^{(m)})\approx$  $\texttt{~{}}\mathbb{P}_{m}(\hat{T}_{1}\mid\hat{T}_{<1},\mathcal{Z}^{(m)})\times\bigl{[}$ (7)
 $\texttt{~{}~{}~{}}\mathbb{P}_{m}(\hat{T}_{2}\mid\hat{T}_{<2},\mathcal{Z}^{(m)})\times\bigl{[}$
@@
 $\texttt{~{}~{}~{}~{}~{}~{}~{}}\mathbb{P}_{m}(\hat{T}_{S}\mid\hat{T}_{<S},\mathcal{Z}^{(m)})$
 $\texttt{~{}~{}~{}~{}~{}~{}~{}}+\sum_{t}\mathbb{P}_{m}(t\mid\hat{T}_{<S},\mathcal{Z}^{(m)})\cdot\mathbb{1}\bigl{(}\{\hat{T}_{<S},t\}\texttt{.prefix}=\{B_{l}\}\bigl{)}$
@@
 $\texttt{~{}~{}~{}}\bigl{]}+\sum_{t}\mathbb{P}_{m}(t\mid\hat{T}_{<2},\mathcal{Z}^{(m)})\cdot\mathbb{1}\bigl{(}\{\hat{T}_{<2},t\}\texttt{.prefix}=\{B_{l}\}\bigl{)}$
 $\texttt{~{}}\bigl{]}+\sum_{t}\mathbb{P}_{m}(t\mid\hat{T}_{<1},\mathcal{Z}^{(m)})\cdot\mathbb{1}\bigl{(}\{\hat{T}_{<1},t\}\texttt{.prefix}=\{B_{l}\}\bigl{)},$
where $t$ represents an alternative token that could follow the sequence $\hat{T}_{<s}$ and meet the criteria of $\{T_{s}\}\texttt{.prefix}=\{B_{l}\}$. Eventually, substituting Equation (7) into (5) can realize generative fusion decoding (GFD).

### v4

# Let’s Fuse Step by Step: A Generative Fusion Decoding Algorithm with LLMs for Robust and Instruction-Aware ASR and OCR

@@ Abstract
We propose “Generative Fusion Decoding” (GFD), a novel shallow fusion framework designed to integrate large language models (LLMs) into cross-modal text recognition systems for automatic speech recognition (ASR) and optical character recognition (OCR). We derive the necessary formulations to enable GFD to operate across mismatched token spaces of different models by calculating likelihood at the byte level, thereby enabling seamless fusion and synchronous progression during the decoding process. GFD is plug-and-play by design, making it readily compatible with various auto-regressive models without the need for any re-training. GFD proves effective for general ASR and OCR tasks through intermediate and frequent interactions with LLMs, surpassing cascaded methods in English and Mandarin benchmarks. In addition, GFD transfers in-context learning abilities of LLMs and allows for adaptive ASR in instruction-aware and long-context settings, yielding significant WER reductions of up to 17.7%.  [^1] Code is available at https://github.com/mtkresearch/generative-fusion-decoding
Let’s Fuse Step by Step: A Generative Fusion Decoding Algorithm with LLMs for Robust and Instruction-Aware ASR and OCR
Chan-Jan Hsu∗1  Yi-Chang Chen∗1  Feng-Ting Liao1 Pei-Chen Ho2Yu-Hsiang Wang2Po-Chun Hsu1Da-shan Shiu1 Equal contribution  1MediaTek Research  2Internship at MediaTek Research {chan.hsu, yi-chang.chen, ft.liao, pochun.hsu, ds.shiu}@mtkresearch.com
@@ 1 Introduction
Figure 1: The GFD integrated framework. The framework aims to integrate pre-trained text-recognition models (ASR/OCR) with LLMs to augment the recognition capabilities. A key challenge in this integration lies in the mismatch between the token spaces of the two model types, which prevents direct fusion. To address this, we derive a formulation (Section 3.1, Equation 7, Equation 8) that enables GFD to compute likelihoods at the byte level, allowing effective fusion during the decoding stage. Here, $\mathcal{Z}^{\texttt{TR}}$ and $\mathcal{Z}^{\texttt{LM}}$ denote the contextual information from the text-recognition model and the LLM, respectively.
@@
Integrating large language models (LLMs) into multi-modal systems has recently emerged as a frontier, significantly advancing applications such as automatic speech recognition (ASR) Radford et al. (2023), visual question answering (VQA) Liu et al. (2023), and reinforcement learning Yang et al. (2023d). Despite their robust capabilities, integrating LLMs with text recognition systems like ASR and OCR poses challenges due to the need for high-quality paired data and extensive training resources. Modern LLMs are trained on trillions of text tokens Hsu et al. (2024); Jiang et al. (2023), far exceeding the data used for end-to-end ASR or OCR models Radford et al. (2023).
Various fusion strategies have been explored in ASR literature, including shallow fusion Chen et al. (2023b); Kannan et al. (2018); Choudhury et al. (2022), late fusion Chen et al. (2024b, a); Xu et al. (2022), mid fusion Radhakrishnan et al. (2023); Liu et al. (2024), and early fusion Fathullah et al. (2024); Chen et al. (2023a). However, these methods face challenges such as discarding the ASR decoder’s denoising abilities Gong et al. (2023) and requiring aligned token spaces. Volatility of model from further training is also a concern when dealing with extensively trained models.
To address these challenges, we introduce a novel shallow fusion framework called “Generative Fusion Decoding” (GFD). GFD operates across mismatched token spaces by calculating likelihood at the byte level, enabling seamless integration of LLMs with text recognition models during the synchronous decoding process (Section 3.1). This plug-and-play framework allows LLMs to correct text recognition errors in real-time (Section 3.2), broadening the exploration space and improving recognition accuracy.
Empirically, GFD is effective on general ASR, especially in challenging scenarios like homophones in Mandarin and code-switching Yang et al. (2023c) (Section 4.2). In addition, GFD transfers long-context awareness and in-context learning Brown et al. (2020b) of LLMs and allows for adaptive ASR. GFD maintains semantic consistency in long-form audio by leveraging transcription history for contextual biasing (Section 4.3). By using controlling prompts such as domain tags, rare words, and explicit instructions, domain sensitivity and instruction awareness is exhibited across various benchmarks (Section 4.4). To the best of our knowledge, this unique aspect of LLM integration has not been reported in prior work Chen et al. (2024b); Hu et al. (2024); Mittal et al. (2024); Hori et al. (2025). Further comparisons with existing methods are discussed in Section 5.
The contributions of this work are summarized as follows:
  • We derive a novel algorithm – GFD, which enables intermediate LLM interaction during the decoding process in text recognition.
  • GFD improves performance on various ASR scenarios and OCR, which is orthogonal to improvements from previous approaches.
  • The robustness of GFD is demonstrated in long-context and instruction-based ASR tasks, which fully utilizes long-range semantic awareness of LLMs. To the best of our knowledge, this has not been reported in prior work with LLM integration.
  • We provide detailed analysis on the performance and the time efficiency of GFD.
@@ 2.1 Model fusion
Training a multi-objective model from scratch is often costly Bapna et al. (2021, 2022); Alayrac et al. (2022); Driess et al. (2023). Consequently, researchers have pivoted towards combining existing models with different modalities to improve accuracy without the prohibitive costs of building new systems from the ground up. Model fusion developed in the field of ASR provides a plausible path for combining existing trained models. The technique have evolved significantly in recent years, encompassing a variety of approaches designed to integrate different models to enhance performance.
Deep fusion integrates models at the level of hidden features, requiring fine-tuning models to fuse deep features Gulcehre et al. (2015). Cross-modal fusion, similar to deep fusion, integrates pre-trained end-to-end ASR model with LLM Radhakrishnan et al. (2023); Yu et al. (2023); Li et al. (2023b) or vision model with LLM Chen et al. (2023a); Liu et al. (2023) via learning a joint representation with large amount of extra paired audio-text or image-text data.
In contrast, shallow fusion or late fusion, often employed in ASR, combines end-to-end ASR models with external language models at the decoding level, improving recognition accuracy without altering the underlying ASR architecture Kannan et al. (2018); Huang et al. (2024); Chen et al. (2024b); Zhang et al. (2023). However, due to the heterogeneous sample spaces of models, the prerequisite of shallow and late fusion requires aligning sample spaces of model distributions, enabled through fine-tuning a projection module Chen et al. (2024b). Late fusion training methods may suffer from modality laziness problem in tasks where uni-modal priors are meaningful Du et al. (2023). Concurrently with our work, step-by-step synchrounous late fusion methods are explored Mittal et al. (2024); Hori et al. (2025). Departing from these efforts, which constrain scoring to specific decoding configurations, our approach addresses the problem from the byte sequence perspective, generalizing the rescoring process to support arbitrary input sequences.
Another line of research integrates LLMs in a cascaded fashion, where the LLM rescores or rewrites based on the N-best hypotheses generated by the first-pass ASR model. While this approach has proven effective in reducing recognition errors Sainath et al. (2019); Hu et al. (2020); Xu et al. (2022), it is limited by the inherently low representation capacity of the N-best list and introduces additional computational latency from the second-pass decoding.
Our newly proposed approach, GFD, operates in the space of homogeneous sequence elements, thereby removing the need for strict token-level alignment. Further, by keeping the model architecture intact, including tokenizers and embedding, we ensure that each individual pre-trained model’s performance on its respective task is preserved and not affected by the instability that may arise from additional training. This property becomes increasingly critical when integrating with large language models (LLMs), which are typically trained on trillions of tokens using carefully refined data curricula and annealed learning rates Dubey et al. (2024).
@@ 2.2 Contextual conditioning
Auto-regressive LLMs have exhibited capabilities in in-context learning Radford et al. , instruction following Ouyang et al. (2022), and knowledge synthesis Liu et al. (2020). Such capabilities have been applied to solving domain adaptation in speech recognition for rare words or out-of-domain context through contextual biasing Choudhury et al. (2022) and prompting fine-tuned models Liao et al. (2023); Yang et al. (2023a); Li et al. (2023b); Yang et al. (2023a). Using GFD, this problem can be addressed by directly leveraging a high-performing LLM through prompting, without the need for additional fine-tuning.
2.3 Mandarin ASR
One of the most significant challenges in developing ASR systems for Mandarin stems from its highly homophonous nature Lee and Chen (1997); Lee (2003); Chen et al. (2022). Unlike English, where there is a larger variety of phonemes and a relatively consistent correspondence between spelling and sound, Mandarin relies on a limited set of tones and syllables to represent thousands of characters. Consequently, Mandarin ASR systems must not only accurately capture the tonal nuances but also analyze the linguistic context to disambiguate these homophones. The integration of LLMs has shown promise in addressing these challenges Chung et al. (2023); Leng et al. (2023); Li et al. (2024), and GFD adopts the same ideology by leveraging the contextual conditioning capabilities of LLMs to enhance Chinese ASR performance.
@@ 3.1 Generative fusion decoding
where $\{T_{s}\}$ represents the sequence of tokens generated by the model, and $\mathcal{Z}$ represents the given context or conditioning information, such as audio for speech recognition models Radford et al. (2023), images for vision-language-models Alayrac et al. (2022), and prompts for typical language models Brown et al. (2020a).
Auto-regressive generation is one approach to realize conditional text generation. In this approach, the auto-regressive model is conditioned on the previously generated tokens to generate the next token sequentially. Therefore, the probability $\log\mathbb{P}(\{T_{s}\}|\mathcal{Z})$ is typically decomposed using the chain rule of probability as follows:
@@
where $T_{s}$ is the token at position $s$ in the sequence, and $T_{<s}$ represents all the tokens preceding position $s$. In real-world applications, it is impracticable to enumerate all possible token sequences, so beam search is typically employed as an approximate strategy to efficiently explore the most likely sequences without exhaustive computation.
In the setting of shallow fusion, multiple models are combined to jointly determine the sequence, as expressed in the following formula, which is a reformulation of Equation (1):
$\{T_{s}^{\texttt{fuse}}\}^{*}$ (3)
$=\underset{\{T_{s}^{\texttt{fuse}}\}}{\mathrm{arg\,max}}\sum_{m}\lambda_{m}\log\mathbb{P}_{m}(\{T_{s}^{(m)}\}=\{T_{s}^{\texttt{fuse}}\},\mathcal{Z}^{(m)})$
where $\{T_{s}^{\texttt{fuse}}\}$ represents the fused sequence of tokens generated by combining the outputs of multiple models, $\lambda_{m}$ is a weighting factor for the $m$-th model, $\mathbb{P}_{m}$ denotes the probability distribution of the $m$-th model and $\mathcal{Z}^{(m)}$ represents the context or conditioning information specific to the $m$-th model.
When the sample spaces of models are the same, Equations (3) and (2) can be combined to realize incremental fusion Chen et al. (2024b). If the models have different sample spaces due to a mismatch in token spaces, there is no simple way to achieve incremental fusion. One alternative method to approximate Equation (3) is to fuse at the level of the fully generated results from each model. Nevertheless, in practice, fusion at the level of fully generated results poses the problem of an enormous search space because different conditioning variables $\mathcal{Z}^{(m)}$ may produce vastly different results.
@@
$\mathcal{M}^{(m)}:\mathbb{P}_{m}(\{T^{(m)}_{s}\},\mathcal{Z}^{(m)})\longrightarrow\mathbb{P}_{m}(\{B_{l}\},\mathcal{Z}^{(m)}),$ (4)
@@
$\{B_{l}^{\texttt{fuse}}\}^{*}$ (5)
$=\underset{\{B_{l}^{\texttt{fuse}}\}}{\mathrm{arg\,max}}\sum_{m}\lambda_{m}\log\mathbb{P}_{m}(\{B_{l}\}=\{B_{l}^{\texttt{fuse}}\},\mathcal{Z}^{(m)}).$
To realize the probability transformation $\mathcal{M}^{(m)}$, we define a mapping from the token-level probabilities to the byte-level probabilities. This mapping takes into account the prefix relationship between the token sequence and the byte sequence. Specifically, we express the byte-level probability $\mathbb{P}_{m}(\{B_{l}\},\mathcal{Z}^{(m)})$ as a sum over all possible token sequences that share a common prefix with the byte sequence $\{B_{l}\}$. The probability of each token sequence is computed as the product of the conditional probabilities of each token given the preceding tokens and the context $\mathcal{Z}^{(m)}$. This relationship is formalized in the following equation:
$\mathbb{P}_{m}(\{B_{l}\},\mathcal{Z}^{(m)})$ (6)
=$=\sum_{\{T^{(m)}_{s}\}}\biggr{[}\prod_{s}\mathbb{P}_{m}(T_{s}^{(m)}\mid T_{<s}^{(m)},\mathcal{Z}^{(m)})\biggr{]}_{\{T^{(m)}_{s}\}}$
×$\times\mathbbm{1}\bigl{(}\{T^{(m)}_{s}\}\texttt{.pref}=\{B_{l}\}\texttt{~{}AND~{}}T^{(m)}_{<s}\texttt{.pref}\neq\{B_{l}\}\bigl{)},$
where .pref is a function that checks whether a sequence $A$ has sequence $B$ as its prefix, and $\mathbbm{1}$ is the indicator function that converts the boolean value of the inner loop to integers ($true\rightarrow 1$, $false\rightarrow 0$). The entire indicator function with two conditions ensures that only the minimal token sequences covering the target byte sequence $\{B_{l}\}$ contribute to the byte-level probability. In Equation (6), the complexity remains high at $\mathcal{O}(V^{S})$, where $V$ represents the vocabulary size of tokens and $S$ is the sequence length, for identifying token sequences that match the specified criteria of the indicator.
@@
This complexity can still be greatly reduced, by eliminating terms with near 0 probability. We posit that the main token sequence and its branching alternatives, as shown in Figure 2, dominate the probability contribution. The main token sequence is produced by applying model tokenization on the byte string and the alternative tokens are essentially look-aheads for potential main tokens that may emerge as the decoding progresses. The significance of the main token sequence is justified by its alignment with the model’s inputs during the original pretraining phase; any other slicing method is penalized in terms of probability due to its lack of representation in the training data. Based on this assumption, we can narrow down the search for token sequences that meet the criteria of ($\bigl{(}\{T^{(m)}_{s}\}\texttt{.pref}=\{B_{l}\}\texttt{~{}AND~{}}T^{(m)}_{<s}\texttt{.pref}\neq\{B_{l}\}\bigl{)}$ to only the main token sequence and its branching alternative tokens. We define the main token sequence as $\{\hat{T}^{(m)}_{s}\}$. Given this simplification, we can approximate the byte-level probability $\mathbb{P}_{m}(\{B_{l}\},\mathcal{Z}^{(m)})$ by considering only the main token sequence and its immediate alternatives that share the same prefix with the byte sequence $\{B_{l}\}$. This approximation significantly reduces the computational complexity to $\mathcal{O}(V\times S)$ and is expressed in the following equation:
$\mathbb{P}_{m,approx}(\{B_{l}\},\mathcal{Z}^{(m)})\approx$ (7)
 $\texttt{~{}}\mathbb{P}_{m}(\hat{T}^{(m)}_{1}\mid\hat{T}^{(m)}_{<1},\mathcal{Z}^{(m)})\times\bigl{[}$
 $\texttt{~{}~{}}\mathbb{P}_{m}(\hat{T}^{(m)}_{2}\mid\hat{T}^{(m)}_{<2},\mathcal{Z}^{(m)})\times\bigl{[}$
@@
 $\texttt{~{}~{}~{}}\mathbb{P}_{m}(\hat{T}^{(m)}_{S}\mid\hat{T}^{(m)}_{<S},\mathcal{Z}^{(m)})$
 $\texttt{~{}~{}~{}}+\sum_{t}\mathbb{P}_{m}(t\mid\hat{T}^{(m)}_{<S},\mathcal{Z}^{(m)})\cdot\mathbbm{1}\bigl{(}\{\hat{T}^{(m)}_{<S},t\}\texttt{.pref}=\{B_{l}\}\bigl{)}$
@@
 $\texttt{~{}~{}}\bigl{]}+\sum_{t}\mathbb{P}_{m}(t\mid\hat{T}^{(m)}_{<2},\mathcal{Z}^{(m)})\cdot\mathbbm{1}\bigl{(}\{\hat{T}^{(m)}_{<2},t\}\texttt{.pref}=\{B_{l}\}\bigl{)}$
 $\texttt{~{}}\bigl{]}+\sum_{t}\mathbb{P}_{m}(t\mid\hat{T}^{(m)}_{<1},\mathcal{Z}^{(m)})\cdot\mathbbm{1}\bigl{(}\{\hat{T}^{(m)}_{<1},t\}\texttt{.pref}=\{B_{l}\}\bigl{)},$
where $t$ represents an alternative token from the token set of the modality $m$. We use $\{\hat{T}^{(m)}_{<s},t\}$ to denote the concatenated sequence of $\hat{T}^{(m)}_{<s}$ and $t$, and this sequence must meet the criteria of leading with $\{B_{l}\}$ to be considered. Eventually, by substituting Equation (7) into $\mathbb{P}_{m}(\{B_{l}\},\mathcal{Z}^{(m)})$ of (5), we have successfully realized generative fusion decoding (GFD). We show that this function is incrementally calculable in Appendix A.1, and thus yields the same time complexity as standard LLM rescoring without branching.